in progress, PBS security breach deflected, systems security breach ongoing with black/red
alert and grid ratio failure, partial recovery. Mplan override command sequences L4/L3 complete, sequence L2 in progress. M
signal field substable, wipeout sequence 8 in progress. Minimum M / DL /
HP / SD requirements met, MDS improving, VTX requirements not met.